# X-LeBench: A Benchmark for Extremely Long Egocentric Video Understanding

### v1

@@ 3.2 Benchmark Tasks
The full-length video is presented to the system only once before querying. The system is required to extract features or frames from the extremely long video and store them in a buffer, responding to query based solely on the stored information. This approach presents a significant challenge to current video understanding systems, which are typically designed to process short videos and thus cannot store information in the presented videos in buffers for future use. Moreover, in real-world scenarios, corresponding answers often recur across multiple time segments as the recording time span increases, so we explicitly annotate the queried time period for each task. This ensures more precise, unambiguous answers, and provides a solution to address the limitations and incompleteness of annotations in existing single-video datasets.
After generating the life-log footage, we selected the corresponding annotations from the Ego4D dataset [12] and adapted them for our novel tasks. Detailed descriptions of each task are provided below. A summary of the number of annotations compiled for each task is presented in Tab. 3.
3.2.1 Temporal Localization
Task Design and Annotation: Temporal localization is critical for enabling virtual personal assistants to handle user queries effectively in egocentric videos. This task, akin to Episodic Memory [12] and Natural Language Video Localization [47, 46], is uniquely challenging for ultra-long videos due to their extensive duration and diverse content. We divide this task into three subtasks based on the type of information being retrieved: objects, people, and moments. For object- and people-related retrieval, we adopt the annotations from the Episodic Memory task (specifically, the Natural Language Query subtask) in Ego4D. Specifically, we retrieve the annotations corresponding to the videos selected by the daily plan chunks (as generated in Sec. 3.1), classify them into topics based on the query templates defined in Ego4D, and remapped the timestamps to align with the virtual time used in the daily plan.
Evaluation Metrics: Following previous work on temporal localization [12, 47, 46], we employed a top-$x$ recall with an intersection over union threshold (“recall@x, tIoU”). $x$ and $t$ are predefined variables that indicate the number of retrieved results we look at and the IoU threshold, respectively. Here, we set $x$ to 5 and $t$ to 0.3.
3.2.2 Summarization
Task Design and Annotation: Summarization is a fundamental task in numerous benchmarks [42, 49, 2], In our context, it provides insights at varying granularities—specific chunks, predefined periods (e.g., morning, afternoon, evening), and holistic summaries. These diverse scopes challenge systems to summarize at multiple summary levels in ultra-long videos. We divide the task into three granularity levels: single-video, multi-video, and holistic. Single-video level focuses on summarizing individual chunks from the daily plan. Multi-video level involves summarizing three predefined time periods: morning, afternoon, and evening. [^1] Morning: before 12 p.m., afternoon: 12 p.m. to 5 p.m., evening: after 5 p.m. Holistic level requires an overall summary of the entire daily footage. To generate summaries at these levels, we employed a hierarchical approach inspired by [16]. Single-video summaries are initially consolidated during the earlier stages of processing (as described in  Sec. 3.1). Using these, we generate multi-video summaries by querying a LLM. The resulting multi-video summaries were then aggregated to produce the holistic summary in a similar manner.
Evaluation Metrics: Although there is an increasing trend to use LLM-based scoring systems for the summarization task (such as in [2]), we found such approaches to be highly unstable and sometimes produce inconsistent scores. Therefore, we instead use traditional summarization metrics such as Rouge-L [22] and the BERT score [48].
3.2.3 Counting
Task Design and Annotation: Counting task tests a system’s ability to accurately retrieve, interpret, and aggregate fine-grained information from multiple parts of the inputs. This task involves counting the occurrences of specific actions within the inputs to reflect user needs for recalling how many times a particular action was performed. Future extensions could involve counting objects or people appearing in the footage.
3.2.4 Ordering
Task Design and Annotation: Ordering task evaluates a model’s capability to recognize temporal relationships between main activities across video records in ultra-long video life logs. Unlike prior studies focusing on fine-grained actions, this task introduces the novel challenge of ordering events based on single-video summaries. Models are tasked with ordering shuffled single-video summaries of entire life logs, posing significant challenges to the system’s temporal understanding ability in long-context inputs. This setup ensures the initial sequence of events is obscured, requiring the model to reconstruct the correct temporal order.
Evaluation Metrics: To evaluate this ordering task, we propose to use a simple accuracy metric. This accuracy is the ratio between the number of correctly predicted order and the total number of summaries.
@@ 4.1 Settings
X-LeBench includes various task types, as outlined in Sec. 3.2, encompassing four categories: temporal localization, summarization, counting, and ordering. These tasks are further divided into eight subtasks, with their respective input-output formats detailed in Tab. 4. Given the interleaved, ultra-long context with chronological attributes in our dataset (Fig. 1), we uniformly construct time-stamped prompts as the long-context input following the format shown in Fig. 4.

### v2

@@ 3.2 Benchmark Tasks
The full-length video is presented to the system only once before querying. The system is required to extract features or frames from the extremely long video and store them in a buffer, responding to query based solely on the stored information. This approach presents a significant challenge to current video understanding systems, which are typically designed to process short videos and thus cannot store information in the presented videos in buffers for future use. Furthermore, as video duration increases, answers to queries may recur across multiple time segments. To address this, we explicitly annotate the queried period for each task, ensuring precise and unambiguous retrieval, while mitigating annotation gaps inherent in single-video datasets. After generating the life-log footage, we adapt corresponding Ego4D annotations for our novel task designs. Tab. 3 summarizes the number of annotations per task.
Method
Data
Category
Object-related
Retrieval (%)
People-related
Retrieval (%)
Moment
Retrieval (%)
Summarization (10)
AVG. Temporal
Localization (%)
AVG.
Summarization (10)
Action
Counting (%)
Ordering
single multi-video holistic
Gemini-1.5-Flash Short 1.67 0.00 10.26 5.12 3.02 4.90 4.20 4.36 13.87 93.70
Medium 1.92 6.90 10.53 4.82 3.12 5.24 5.88 4.46 16.74 44.66
Long 1.75 11.11 8.33 4.57 3.44 5.33 6.45 4.43 16.76 24.40
Socratic Models Short 8.33 25.00 17.95 6.80 3.69 6.47 14.28 5.67 9.53 85.59
Medium 5.77 20.69 18.42 6.64 4.09 6.41 13.45 6.04 14.04 55.63
Long 12.28 22.22 10.83 5.82 4.37 6.15 11.83 5.61 19.85 23.33
Retrieve - Socratic Short 8.33 25.00 17.95 6.88 3.79 6.53 14.28 5.75 9.92 87.67
Medium 9.62 17.24 7.89 7.01 4.30 6.56 10.92 6.36 9.38 54.55
Long 10.53 66.67 14.17 6.84 4.69 6.48 15.59 6.48 9.77 24.17
Gemini-1.5-Flash All 1.77 5.17 9.14 4.73 3.20 5.16 5.66 4.43 16.44 41.96
Socratic Models All 8.87 22.41 13.70 6.22 4.06 6.34 12.97 5.76 17.15 42.61
Retrieve - Socratic All 9.47 27.59 13.71 6.90 4.28 6.52 13.92 6.30 9.67 43.01
Table 5: Evaluation results on X-LeBench. Including temporal localization tasks (object-related, people-related and moment retrieval), summarization tasks (single-video, multi-video and holistic level), action counting and summary ordering task. AVG. temporal localization and summarization is the average performance of all temporal localization tasks and summarization tasks, respectively.
Figure 4: Performance comparison on X-LeBench. (a): Radar chart showing the overall performance of each method across all tasks. For fair comparison across tasks, scores for summarization are scaled by a factor of 10 (maximum 100). (b): The performance comparison across different data categories for different tasks.
@@ 4.1 Settings
X-LeBench includes various task types, as outlined in Sec. 3.2, encompassing four categories: temporal localization, summarization, counting, and ordering. These tasks are further divided into eight subtasks, with their respective input-output formats detailed in Tab. 4. Given the interleaved, ultra-long chronological context (Fig. 1), we uniformly construct time-stamped prompts as the long-context input following the format shown in Fig. 5.
To reflect real-world usability, tasks require free-form textual outputs instead of closed-set multiple-choice answers. This approach increases practical usability while significantly raising the complexity of tasks, as it demands more nuanced understanding and reasoning. To ensure the evaluability of outputs, the prompts provide extended context along with task-specific instructions and output format requirements for each task type, standardizing the expected responses. For temporal localization and summarization, each query is independently evaluated, while for counting and ordering, we aggregate the query contents of each data entry to action or order lists for a unified assessment. Notably, due to computational cost and time associated with analyzing ultra-long contexts, along with their overall poor performance and limited analytical value—we test only a limited set on temporal localization.
Overall, X-LeBench presents significant challenges for current video understanding systems, which are often limited by short context handling, weak temporal reasoning, and poor long-horizon memory retrieval. To evaluate performance under these constraints, we test three representative multi-modal approaches to handle ultra-long contexts:
(1) Gemini-1.5-Flash team2024gemini: A native multi-modal model excels in handling ultra-long contexts and is trained jointly on multi-modal data. It is used in an end-to-end manner by uniformly sampling 1200 frames per data, with a temperature of 0.1.
(2) Socratic Models zeng2022socraticmodels: Most advanced multi-modal systems cannot process ultra-long videos. Inspired by related works chandrasegaran2024hourvideo; zeng2022socraticmodels, we divide videos into 30-second segments and generate captions for these segments using the Qwen-VL-7B model Qwen2-VL, captioning segments at 1 frame per second and resolution settings of $560\times 420$. The captions are then aggregated with timestamps as life-log records, forming the input for executing long-context understanding tasks. The Gemini-1.5-Flash model is used for performing tasks on these textual life log records, with the same parameters as in method (1).
(3) Retrieve-Socratic: Considering the inherent temporal nature of our dataset, we propose an enhanced version of the Socratic Models approach. This method incorporates a rule-based filtering mechanism to extract only task-relevant time segments, reducing irrelevant temporal data. By narrowing the contextual focus, this approach aims to improve efficiency and relevance in long-context understanding. The extracted contextual information is then processed using the same Socratic Models’ pipeline and parameter settings as in method (2).
We also evaluate the LongVU shen2024longvu, an open-source method for long-form video understanding. Detailed results are in Appendix E.
4.2 Results
As shown in Fig. 4 and Tab. 5, we evaluate the performance of three methods across short, medium, long, and all data categories. The “all” category represents an aggregated evaluation of the former 3 data types, offering a more comprehensive and holistic assessment. The evaluation metrics for each task are detailed in Appendix C.
Overall Performance. As shown in last three rows of Tab. 5 and Fig. 4 (a), the Retrieve-Socratic method demonstrates superior performance in most tasks. Due to input token limitations, longer videos require larger frame sampling intervals, resulting in increased information loss. Consequently, both Socratic and Retrieve-Socratic methods significantly outperform Gemini-1.5-Flash on temporal localization and summarization tasks. Specifically, Retrieve-Socratic achieves an average improvement of 8.26% in recall for temporal localization and 1.87 points higher scores in summarization. The underperformance of Gemini-1.5-Flash in temporal localization reflects its difficulty in fine-grained temporal reasoning under long-context constraints. Additionally, Retrieve-Socratic further improves upon the basic Socratic approach by narrowing the temporal context, leading to notable gains in both temporal localization and summarization tasks. In holistic summarization and ordering, where the queried information spans the entire video, the performance of Socratic and Retrieve-Socratic methods remains comparable, demonstrating their robustness in global reasoning.
Interestingly, our analysis of the specific results indicates that the Retrieve-Socratic method tends to output 0 more frequently in the counting task. This leads to noticeably poorer performance compared to other methods, suggesting that excessive context reduction may introduce unintended biases in certain tasks.
Impact of Data Types. As illustrated in Fig. 4 (b) and Tab. 5, the performance trends across data types are not uniform. These trends may be influenced by varying task-specific objectives and understanding burden imposed by different data types on each method. Overall, both Socratic and Retrieve-Socratic demonstrate more stable performance trends across data types compared to Gemini-1.5-Flash. This suggests that the processing capability of textual information is more robust when dealing with such long-context information. For the ordering task, all methods achieve strong performance on short videos (accuracy exceeding 85%). However, as video length increases, performance across all methods declines significantly, with accuracy dropping below 25% for long data.
Impact of Temporal Structure. To explore how the complexity of temporal structures impacts the difficulty of core tasks involving temporal understanding, we conducted experiments on a “shorter” split of our dataset, where each video life log consists of only two videos. This setup enables controlled analysis of temporal reasoning under reduced structural complexity. The comparative results with the “short” split are presented in Tab. 6.
Method
Single-video
Sum. (10)
Multi-video
Sum. (10)
Ordering
AVG.
Temporal Loc. (%)
Gemini-1.5-Flash
5.52
($\uparrow$ 0.40)
4.20
($\uparrow$ 1.18)
97.20
($\uparrow$ 3.50)
0.84
($\downarrow$ 3.36)
Socratic Models
6.92
($\uparrow$ 0.12)
4.92
($\uparrow$ 1.23)
97.22
($\uparrow$ 11.63)
14.29
($\uparrow$ 0.01)
Retrieve-Socratic
7.00
($\uparrow$ 0.12)
5.89
($\uparrow$ 2.10)
96.88
($\uparrow$ 9.21)
17.65
($\uparrow$ 3.37)
Table 6: Performance comparison on the “shorter” split. Arrows indicate relative differences compared to the “short” split.
We observe a clear performance increase on the “shorter” split across summarization and ordering tasks, indicating that as temporal structure becomes more complex, models suffer substantial degradation even in relatively coarse-grained temporal reasoning. In addition, performance on temporal localization tasks remains consistently poor across all settings—Gemini-1.5-Flash’s result even declines on shorter inputs. This highlights that fine-grained temporal grounding is an intrinsic bottleneck, independent of sequence length or structure. The task requires pinpointing events at the granularity of a few seconds within videos spanning tens of minutes, a challenge exacerbated by token budget limits and information bottlenecks. General-purpose models are not well-optimized for this setting, and while recent methods show progress on shorter videos, performance on extended sequences remains severely constrained.
Remarks. (1) Model refusal rate: We observe a notable refusal rate from Gemini-1.5-Flash due to its built-in information security constraints, particularly when directly processing long-form video input. The refusal rates increase with input length, reaching 20.14%, 27.08%, and 30.56% for short, medium, and long videos, respectively. Both the Socratic and Retrieve-Socratic methods, which convert video content into textual descriptions before prompting the model, achieve a 0% refusal rate. LongVU does not exhibit any refusal behavior. (2) Temporal reasoning failure: In the multi-video summarization task, a significant portion of responses are invalid, even when timestamps are included in input context, and the query time is explicitly specified in the question (e.g., “12:00 to 17:00”). Typical invalid responses include statements like “There is no information available within the specified time period”, indicating a breakdown in the temporal understanding of the model. The failure rates of the multi-video summarization task are 55.28%, 51.89%, and 57.03% for Gemini-1.5-Flash, Socratic Models and Retrieve-Socratic, respectively. LongVU demonstrates almost no temporal awareness, it consistently disregards the specified time window and instead summarizes the entire video content. Also, in the ordering task, its performance on short, medium, and long data drops drastically to 37.5%, 1.7%, and 0.69%, respectively. Manual review of the outputs reveals that even on short data, LongVU frequently returns meaningless outputs like “0, 1, 2, 3” without establishing any meaningful temporal structure. Therefore, we only included LongVU in tasks that require global summaries or coarse-grained time reasoning (i.e., holistic summarization and ordering). Additional results and analysis are in the Appendix E.
4.3 Summary of Key Findings
Temporal Reasoning: The Core Bottleneck. Evaluated models struggle with temporal reasoning, showing high failure rates in multi-video summarization and localization. This highlights the need for models with improved temporal alignment and contextual memory for long sequences.
Textual Representations: A Path Towards Scalable Understanding. Our findings suggest that structured textual representations serve as a highly effective and scalable intermediate for ultra-long video understanding. As demonstrated by the Socratic and Retrieve-Socratic methods, converting raw long video inputs into textual forms significantly reduces model refusal rates and consistently enables more stable and improved performance. This strategy not only effectively alleviates the severe token limitations of current multi-modal models but also dramatically reduces computational overhead. This makes it a practical and highly scalable alternative for long-form video understanding, especially in scenarios where direct processing of extended video sequences remains challenging.
Context Filtering: Balancing Efficiency and Completeness. The Retrieve-Socratic approach leverages rule-based temporal filtering to retain only task-relevant context, resulting in superior performance in temporal localization and summarization tasks. However, its relatively weaker performance in counting task suggests that aggressive pruning may omit essential contextual cues. Going forward, this insight underscores the need to design more adaptive retrieval-augmented generation (RAG) systems capable of balancing context filtering with comprehensive long-term information retention in ultra-long video understanding.
5 Conclusion
We introduce X-LeBench, the first benchmark for ultra-long egocentric video understanding, featuring a customizable life-logging simulation pipeline. By curating 432 video life logs from Ego4D, categorized into short, medium, and long durations. X-LeBench provides a diverse, structured dataset for long-form video analysis. It includes tasks in temporal localization, summarization, counting, and ordering, offering a rigorous evaluation framework. We hope to advance research in long-form egocentric video processing, fostering the development of more robust and temporally aware AI models.
6 Limitations
Despite our efforts to construct an extremely long video dataset with contextually coherent activity contents, limitations remain due to the scarcity and insufficient diversity of available data. For details of videos, contextual inconsistencies exist within the dataset, such as different kitchen or office settings within the same data. However, since our focus is on long-term activity content, these discrepancies are beyond the scope of this work. In addition, while MBTI is employed in our simulation pipeline to introduce behavioral diversity, we acknowledge that it is a heuristic framework with limited scientific validity. Its inclusion serves a pragmatic purpose—to support structured variation in persona-driven daily plans—and is not intended to imply psychological rigor or generalizable personality modelling.
Future work could involve expanding both the modalities and time spans of collected data through diverse devices and sources to enhance a comprehensive understanding of long-form human activity, supporting the development of more personalized tasks that emphasize long-term memory or preference-centric reasoning and predictive capabilities.
Appendix A Persona Profile
To simulate diverse egocentric behaviors, we generated 144 unique persona profiles by combining 16 MBTI personality types with 9 different location settings. Each profile serves as input for our dataset creation. Tab. 7 presents a detailed example of a persona profile, illustrating its structure and content.
Field Description
Location UK
Personality Traits MBTI Type: INFP Character Traits: Idealistic and creative, often daydreaming about a better world Highly empathetic and compassionate, with a strong sense of personal values Prefers deep, meaningful connections over a wide social circle Flexible and adaptable, but can be disorganized and easily overwhelmed by details Values authenticity and is highly attuned to the feelings of others
Lifestyle C wakes up around 7:30 AM and starts the day with meditation or journaling. Works from home as a freelance writer from 9:00 AM to 5:00 PM, with breaks for lunch and short walks. Evenings are spent reading, writing poetry, or engaging in creative projects, and bedtime is around 11:00 PM.
Hobbies Writing poetry and short stories Reading literature and philosophy Practicing mindfulness and meditation Volunteering at local community centers Exploring nature and taking long walks in the countryside
Daily Agenda Wake up at 7:30 AM Meditation or journaling from 8:00 AM to 8:30 AM Start work at 9:00 AM Lunch break from 1:00 PM to 2:00 PM End work at 5:00 PM Evening walk or light exercise from 6:00 PM to 7:00 PM Dinner and relaxation from 7:00 PM to 9:00 PM Reading or creative activities from 9:00 PM to 10:30 PM Go to bed at 11:00 PM
Daily Plan Chunks 07:30–08:00: Wake up and stretch 08:05–08:35: Meditation or journaling 08:40–09:10: Prepare a healthy breakfast and make tea 09:15–09:45: Eat breakfast, clean the kitchen, and set up the workspace 09:50–10:50: Start work as a freelance writer 10:55–11:00: Continue working on writing projects 11:05–11:20: Take a short break to stretch and hydrate 11:25–13:00: Continue working on writing projects 13:05–14:00: Lunch break and light reading 14:05–16:00: Resume work and handle emails and client communications 16:05–17:00: Continue handling emails and client communications 17:05–17:35: End work and tidy up the workspace 17:40–18:40: Evening walk or light exercise in the park 18:45–20:00: Prepare and eat dinner, relax with some music 20:05–21:35: Engage in creative activities (writing poetry, reading literature, or working on a personal project) 21:40–22:30: Wind down with light reading or calming music
Table 7: Persona Profile: Tabular representation of an INFP individual’s daily life, lifestyle, hobbies, and activities.
Appendix B Coarse-to-Fine Video Matching and Iterative Optimization
As described in Sec. 3.1, the life-log simulation pipeline generates multiple daily activity chunks, which are then matched with videos from the video selection library to construct the dataset. Stage 3 follows a coarse-to-fine matching approach based on Equation 1 with the implementation details outlined below:
B.1 Coarse Matching ($M_{v}(\cdot)$)
For each activity chunk $C$, key information is extracted, including $C_{t}$ (time period), $C_{scene}$ (scene type), and $C_{scenario}$ (scenario). The $C_{scene}$ and $C_{scenario}$ attributes are inferred using GPT-4o based on the chunk’s textual description $C_{desc}$. Candidate videos are retrieved from the video selection library where the resource region matches the persona’s predefined location. A video is included in the candidate set $M_{v}(\cdot)$ when its time and scene type match the $C_{t}$ and $C_{scene}$, with an overlap between its scenarios and $C_{scenario}$ larger than 0.33 (i.e., the IoU of their scenarios). Note that, if no matching videos are found, the criteria are relaxed by lowering the overlap threshold to 0.2 or removing the regional constraint. This process produces the initial candidate set $M_{v}(\cdot)$.
B.2 Fine Matching ($S(\cdot)$)
From the candidate set $M_{v}(\cdot)$, a finer selection is made by computing the sentence similarity $S(\cdot)$ between the chunk’s textual description $C_{desc}$ and the textual summaries of the candidate videos. The two videos with the highest similarity scores are identified, and one is randomly selected to increase dataset diversity while maintaining relevance.
B.3 Iterative Optimization
Once a matching video is selected, its metadata is incorporated into the persona’s memory as a record of completed activities. This updated memory, along with the persona’s predefined attributes, is provided to GPT-4o to perform a reflection process, enabling the model to refine and adjust the upcoming plan chunks to maintain logical coherence. The next plan chunk is then processed using the updated persona memory, and coarse-to-fine matching is repeated iteratively. This process continues until all plan chunks are matched with corresponding videos, ensuring contextually coherent and realistic life-log simulations.
Appendix C Task Details
As described in Sec. 3.2, X-LeBench provides a comprehensive benchmark for daily activity-related tasks, comprising four major categories with eight subtasks. Here, we detail the task settings and the prompt templates used during evaluation. The prompt templates and example outputs are illustrated in Tab. 4 and Fig. 5.